OpenStudy (anonymous):

find dy for y=e^((square root of x)-1)

OpenStudy (anonymous):

what's the derivative of e^x?

OpenStudy (anonymous):

e^x

OpenStudy (anonymous):

In solving an equation for x, 3 is subtracted from one side of the equation. What needs to be done to the other side of the equation? A. add 3 B. divide by 3 C. subtract 3 D. nothing

OpenStudy (anonymous):

right but in this question you're given a function instead of x, so which rule would you apply?

OpenStudy (anonymous):

In solving an equation for x, 3 is subtracted from one side of the equation. What needs to be done to the other side of the equation? A. add 3 B. divide by 3 C. subtract 3 D. nothing

OpenStudy (anonymous):

jp12sky, you can post your question where it says ask a question

OpenStudy (anonymous):

In solving an equation for x, 3 is subtracted from one side of the equation. What needs to be done to the other side of the equation? A. add 3 B. divide by 3 C. subtract 3 D. nothing

OpenStudy (anonymous):

you use the chain rule right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

can you do it now?

OpenStudy (anonymous):

In solving an equation for x, 3 is subtracted from one side of the equation. What needs to be done to the other side of the equation? A. add 3 B. divide by 3 C. subtract 3 D. nothing

OpenStudy (anonymous):

3 would be subracted from other side too jp12sky

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

\[f(x)=e^{\sqrt{x}-1}\] like that?

OpenStudy (anonymous):

okay i didnt read the question right but yes i know how to do it, you pretty much find the derivative.

OpenStudy (anonymous):

right so e^(anything) = e^(anything) and then you multiply it with the derivative of that anything